50 ml of hard water consumed 15 ml of 0.01 EDTA before boiling.Calculate the degree of total hardness,temp and perm hardness
 

Grade:12th pass

2 Answers

Arun
4 years ago
50ml of water sample=15ml of 0.01M EDTA
= ml of 0.01 EDTA=300ml of 0.01M EDTA
=2*300ml of 0.01 N EDTA
=600 ml or 0.6 L of 0.01 eq. of CaCO3
=0.6*0.01*50g CaCO3 eq.
Then the total hardness=0.30g or 300mg of CaCO3 eq=300mg/L or ppm
now 50 ml of boiled water=5 ml of 0.01 M EDTA
1000ml of boiled water= ml of 0.01 M EDTA
=100 mL of 0.01 M EDTA
=200ml or 0.2 L of 0.01 N EDTA
=0.2*0.01*50 g of CaCO3 eq
=0.1 g or 100mg of CaCO3 eq
permanent hardness=100mg/L or ppm
Temporary hardness=300-100 = 200 ppm
Tashu
one year ago
Calculate the hardness of water sample is 20 ml required 30 mld A 10 ml of standard ccl2 solution the equivalent to 300 mg of caco through pers 200 ml required 20 ml of the same Adda solution on the other hand 50 ml of the boil water sample when they divided against consumed 10 ml of solutions of calculate total permanent and Temporary hardness